OpenStudy (keana):

explain the two ways to simplify √147

OpenStudy (anonymous):

factor out 147 into 3 x 49. 49 is a perfect square

OpenStudy (anonymous):

\[\sqrt{147} = \sqrt{49*2}\] Can you do it from here?

OpenStudy (anonymous):

Crap, sorry, replace the 2 with a 3.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

\[\sqrt{149} = 49 \times 3 = 7\sqrt{3} \] So this is simplified .